OpenStudy (anonymous):

how do you solve the system x+y=5 and x+y=3

geerky42 (geerky42):

You can't. Because \(5\neq3\)

geerky42 (geerky42):

If x+y = 5 and x+y = 3, then we have 5 = x+y = 3 yield 5=3, which is false equation

geerky42 (geerky42):

So there is no solution